OpenStudy (anonymous):

PLSSSSSSS HELP Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

OpenStudy (anonymous):

Did you look at any of those links I sent you? (teach a man to fish = lifetime independence)

sam (.sam.):

First multiply out the brackets

OpenStudy (anonymous):

3x+1 = -2x -2 +6 ??

sam (.sam.):

yes then put all the coefficients with variables into one side

OpenStudy (s):

3(x+1)=-2(x-10+6 3x+3=-2x+2+6 3x+2x=2+6-3 5x=5 x=5/5 x=1

sam (.sam.):

3x+3 = -2x -2 +6

sam (.sam.):

5x=5 x=1

OpenStudy (anonymous):

how about this one Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

sam (.sam.):

x^2+1=0 x^2=-1

OpenStudy (anonymous):

can you explain how you did the steps please

sam (.sam.):

\[x^2+1=0 \] Substract -1 to both sides, \[x^2=-1 \] square root both sides \[x=\sqrt{-1}\] Since square root of -1 contains imaginary number only, there is no real number solution to \[\sqrt{-1}\]